Sine inverse root3-1 by 2root2Please prove this by whole steps

Grade:12

1 Answers

Rahul
5 years ago
Sin^-1(√3-1/2√2) = xThe value of (√3-1/2√2) is 0.2588Sin^-1(0.2588) =xπ/12=x[X=π/12]CCCXXXXxXX